2 A PaRer by L.F. Richaideon Pu blished in the Joumal of the Royal suatistical Society orol忙cd thè distribution af wars over time ad Concluded that wars that begin ing a colender year hod a poisoan distriaion wth parameter2-o.0t, what is the Probability tha there will be at least two wars daring a given Clender year
Answer #1

Let X is a radnom vaiable shows the number of wars during a calender year. Here X has Poisson distribution with following pdf:

P(X=x)=\frac{e^{-0.07}0.07^{x}}{x!},x=0,1,2,..

The probability that there will be at least 2 wars during a calender year is

P(X\geq 2)=1-P(X\leq 1)=1-\sum_{x=0}^{1}\frac{e^{-0.07}0.07^{x}}{x!}=1-0.99766=0.00234